40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Coating Ring-down Measurement Lab elog  Not logged in ELOG logo
Message ID: 103     Entry time: Tue Aug 23 09:21:25 2016
Author: Gabriele 
Type: General 
Category: General 
Subject: Status of Mark Optics disks 

We initially received 20 disks (75 mm diameter, 1 mm thickness) from Mark Optics. Here's their status as of today

  • MO-01 was annelaed and measured at LMA. It was then taken back to Mark Optics by Julie Houser at the time of her visit
  • MO-02 was initially used to test the "dirty" measurement apparatus, then it was installed into the new chamber and measured on Saturday August 20th
    Then this disk was used as a guinea pig for the laser edge polishing, and it's now in a pretty bad shape
  • MO-03 was installed into the new chamber and measured a couple of times on Sunday and Monday. Then we laser polished the edges, with good results, although there are small damages on the surface. It has been installed back into the chamber and a couple of rind-down measurement carried out on Monday and Tuesday.
  • MO-04 is still untouched

All the other disks have been sent back to Mark Optics to grind out flats. 

ELOG V3.1.3-